# Volumetric Semantically Consistent 3D Panoptic Mapping

### v1

@@ Abstract
We introduce an online 2D-to-3D semantic instance mapping algorithm aimed at generating comprehensive, accurate, and efficient semantic 3D maps suitable for autonomous agents in unstructured environments. The proposed approach is based on a Voxel-TSDF representation used in recent algorithms. It introduces novel ways of integrating semantic prediction confidence during mapping, producing semantic and instance-consistent 3D regions. Further improvements are achieved by graph optimization-based semantic labeling and instance refinement. The proposed method achieves accuracy superior to the state of the art on public large-scale datasets, improving on a number of widely used metrics. We also highlight a downfall in the evaluation of recent studies: using the ground truth trajectory as input instead of a SLAM-estimated one substantially affects the accuracy, creating a large gap between the reported results and the actual performance on real-world data. The code is available at: https://github.com/y9miao/ConsistentPanopticSLAM.
@@ I INTRODUCTION
3D semantic-instance mapping is the problem of simultaneously detecting and segmenting objects with their semantic and instance labels to generate a 3D semantic map of the surrounding environment. There are two popular approaches to this problem: 3D-to-3D, which analyzes input 3D point clouds – in this scenario the environment is previously reconstructed; and 2D-to-3D, which analyzes a set of 2D images, further projecting and aggregating predictions on the 3D map – here reconstruction and semantic mapping is done synchronously. In practice, 3D-to-3D approaches often suffer from a lack of large-scale training data and speed issues, i.e., 3D data is more complex than 2D, which makes processing and analyzing 3D point clouds computationally intensive. 2D-to-3D algorithms leverage state-of-the-art 2D semantic-instance segmentation algorithms hence resolving some of the issues with 3D-to-3D approaches. However, having only 2D observations of the 3D instances, such methods often produce a noisy segmentation of the environment due to the limited perspective of 2D images and the noise in semantic detections per frame and camera calibration.
@@ II RELATED WORK
Among the representative techniques are Voxblox++ [1] and PanopticFusion [14]. Voxblox++ [1] employs both semantic instance and geometric segmentation, converting 2D frames into instance-labeled 3D segments. The process involves the incremental registration of geometric segments detected in depth frames onto global 3D counterparts, subsequently integrating semantic and instance labels within these global segments. Our method is inspired by [1], but it offers substantial enhancements, particularly in terms of semantic accuracy. In [14], panoptic segmentation [15] is applied to extract semantic instances for “things” and semantic labels for collective objects (“stuff”), which help improve the quality of semantic-instance segmentation. SceneGraphFusion [16] builds a scene graph with geometric segments and then applies a graph neural network to predict semantic and instance labels. However, SceneGraphFusion utilizes depth images synthetically rendered directly from a reconstructed 3D mesh, hence using ground truth data. The recent INS-CONV [17] combines the advantages of 2D-3D and 3D-3D procedures. While it builds the 3D model of the environment incrementally from RGB-D frames, it performs segmentation directly in the 3D space. This, however, comes at the cost of requiring 3D ground truth annotations for training.
@@ III SEMANTIC-INSTANCE MAPPING
The proposed incremental 2D-to-3D instance and semantic mapping approach, shown in Fig. 1, takes a sequence of RGB-D frames and poses estimated by SLAM pipeline, such as [22], as input, and goes through a six-step procedure for each RGB-D frame: (i) 2D panoptic-geometric segmentation (Sec. III-A); (ii) 3D superpoint segmentation (III-B); (iii) superpoint graph construction and update (III-C). The next steps run only when corresponding information is queried: (iv) semantic regularization (III-C); (v) instance refinement (III-D); and (vi) mesh generation as in [1].
@@
Panoptic-geometric segmentation (Fig. 1 III-A), is applied on each RGB-D frame $F$ based on the assumption [18]:
@@
The surface of each real-world object consists of single or multiple neighboring convex surfaces which share the same semantic-instance label.
First, panoptic segmentation [23] on the RGB image provides a set of semantic instance masks and geometric segmentation [18] on the depth image extracts a set of convex segments. Intersections between panoptic and geometric segmentation masks are used to generate a set of surface masks $\mathcal{F}_{s}=\{s_{i}\}$ per frame $F$. In this way, each surface mask $s_{i}$ is frame-wise uniquely associated with a panoptic semantic instance $o_{i}$ and a geometric segment, thus being less prone to geometric and semantic under-segmentation.
Each surface $s_{i}$ is assigned with instance label $o_{i}$, semantic category $C(o_{i})$, and panoptic confidence defined as:
@@
where $C_{\text{Th}}$ and $C_{\text{St}}$ are the sets of “thing” and “stuff” categories [23]. $\text{Score}(o_{i})$ is the confidence score for instance $o_{i}$ from [23]. Given the camera pose and depth image, the point cloud $\text{Pcl}(s_{i})$ of $s_{i}$ is projected to a unified Voxel-TSDF structure, that is incrementally built.
III-B Semantically Consistent Superpoint Segmentation
In addition to the above, each surface $s_{i}\in\mathcal{F}_{s}$ is also associated with a superpoint label $L_{S}(s_{i})$. Next, we explain what is a superpoint and how to compute the association between a surface and a superpoint.
Based on Assumption 1, we define a “superpoint” $S$ as the set of neighboring convex surfaces stemming from the same 3D object. Each superpoint has a unique identifier, or else label, $L_{S}$. This leads to the following assumption:
@@ III-B Semantically Consistent Superpoint Segmentation
Each real-world object consists of single or multiple spatially connected superpoints with the same semantic-instance labels.
Based on Assumptions 1 and 2, we interpret surface $s_{i}\in\mathcal{F}_{s}$ as a partial observation of a superpoint $S$ in the global 3D space. Each surface $s_{i}$ is assigned to a superpoint by calculating the overlap $\Pi_{s}$ between $s_{i}$ and candidate superpoints in the 3D space (i.e., the number of overlapping voxels in the Voxel-TSDF structure). Specifically, for each point $p_{i}$ of $\text{Pcl}(s_{i})$, we query the superpoint label $L_{S}$ in voxel $V(p_{i})$ and then update the overlap count incrementally as $\Pi_{s}(L_{S}(V(p_{i})),s_{i}):=\Pi_{s}(L_{S}(V(p_{i})),s_{i})+1$. Surface $s_{i}$ is assigned to superpoint $L_{S}$ if the overlap is large enough, controlled by a hyper-parameter. Otherwise, a new label $L_{S_{\text{new}}}$ is created for $s_{i}$.
To ensure that assumption 2 is met, we maintain $\Pi_{V}(L_{S},V_{i})$, the count of occurrences of superpoints $L_{S}$ in a voxel $V_{i}$. The final superpoint label $L_{S}^{*}$ for $V_{i}$ is
$L_{S}^{*}(V_{i})=\arg\max_{L_{S}}\Pi_{V}(L_{S},V_{i}).$ (2)
i.e., a voxel is assigned to the superpoint with the most votes. We use the above equation when incorporating a new surface $s_{i}$ in the global map and querying the current superpoint label $L_{S}$ in voxel $V(p_{i})$.
To address the over-segmentation of superpoints, a pairwise (voxel) overlap score $M(L_{S_{1}},L_{S_{2}})$ is maintained and updated for each surface $s_{i}$ as follows:
$\small M(L_{S_{1}},L_{S_{2}})=\sum_{t=0}^{n}\sum_{s_{i}\in\mathcal{F}_{s}}\llbracket S_{1}\in L_{N}(s_{i})\wedge S_{2}\in L_{N}(s_{i})\rrbracket,$ (3)
@@
  1. 1.
    Spatially connected: $M(L_{S_{1}},L_{S_{2}})>\theta_{\text{merge}}$.
  2. 2.
    Semantically consistent: $C_{S}(L_{S_{1}})$ equals to $C_{S}(L_{S_{2}})$ or at least one of $C_{S}(L_{S_{1}}),C_{S}(L_{S_{2}})$ equals to $C_{0}$, where $C_{S}$ is the semantic label of a superpoint, determined by Eq. 4 and $C_{0}$ is the background semantic label.
Figure 2: Superpoint segmentation without (left) and with (right) semantic consistency as proposed in Section III-B
Considering semantic consistency helps ensure that Assumption 2 holds and, thus, achieves higher superpoint segmentation accuracy than [1, 21]. An example is shown in Fig. 2.
@@ III-C Semantic Segmentation via Super-Point Graph
To address that 2D instance labels are often inconsistent across frames since inference happens frame-wise, prior work [1, 21] apply a data association step to find correspondences of 2D instances and 3D instance labels. Essentially, this process counts observations of semantic-instance labels and assigns superpoints to the labels with the maximum count. However, this method has its pitfalls. It simplistically assigns an instance to the superpoint that has the greatest overlap. As a result, it tends to miss fine details, leading to under-segmentation. For instance, smaller objects near larger ones might be misidentified as a part of the bigger entity (like mistaking a mug as a part of a table). Conversely, it is prone to over-segmentation, especially when there is minimal overlap between the frame-wise segments and superpoints, often due to inaccuracies in pose or depth measurements. This often leads to superfluous instance labels in practice.
To tackle the aforementioned challenges, we leverage a superpoint graph to aggregate frame-wise semantic-instance-spatial data, which aids in determining the semantic-instance label assigned to a superpoint. Let us construct a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with superpoints as vertices ($L_{S}\in\mathcal{V}$). We connect two vertices by edge $\epsilon^{C}=(L_{S_{1}},L_{S_{2}})\in\mathcal{E}$ if the spatial-panoptic confidence $P_{\epsilon}^{C}(L_{S_{1}},L_{S_{2}})>0$. Confidence $P_{\epsilon}^{C}(L_{S_{1}},L_{S_{2}})$ denotes how likely two superpoints $L_{S_{1}},L_{S_{2}}$ share the same instance label with semantic class $C$ and are spatially connected. Each node has its semantic confidence $P_{\nu}^{C}(L_{S})$ representing how likely $L_{S}$ belongs to semantic $C$.
Confidence calculation. Here, we describe how confidence $P_{\nu}^{C}$ and $P_{\epsilon}^{C}$ is calculated. For frame $F$, we ray-cast the pixels of each instance $o_{i}$ into the 3D superpoint space. The confidence being assigned to class $C(o_{i})$ in superpoint $S$ comes from aggregating the voxels of $S$ that intersect the rays. Precisely, $P_{\nu}^{C}(L_{S})=P_{\nu}^{C}(L_{S})+P_{o_{i}}\cdot N^{S}_{o_{i}}$, $N^{S}_{o_{i}}=\sum_{V\in\text{voxels}_{L_{S}}}\llbracket\text{intersect}(V,\text{rays}(o_{i}))\rrbracket$. If two superpoints are hit by rays of $o_{i}$, pairwise confidence is updated: $P_{\epsilon}^{C(o_{i})}(L_{S_{1}},L_{S_{2}})=P_{\epsilon}^{C(o_{i})}(L_{S_{1}},L_{S_{2}})+P_{o_{i}}\cdot P_{spatial}\cdot min(N^{S_{1}}_{o_{i}},N^{S_{2}}_{o_{i}})$, $P_{spatial}$ is the spatial confidence determined by spatial distance between $L_{S_{1}}$ and $L_{S_{2}}$. The initial guess of semantic labels to superpoints is assigned and updated during incremental mapping:
$C(L_{S})=\max_{C}P_{\nu}^{C}(L_{S}),$ (4)
Graph optimization. However, such assignment is based only on individual observations of each superpoint. To further optimize this assignment, we propose a graph-based approach employing the $\alpha-\beta$ swap algorithm [24].
Given graph $\mathcal{G}$ constructed as described previously, we apply the $\alpha-\beta$ swap algorithm, minimizing energy
$\small E(C)=\sum_{L_{S}\in\mathcal{V}}\psi_{u}(C(L_{S}))+\sum_{(L_{S_{1}},L_{S_{2}})\in\mathcal{E}}\psi_{p}(C(L_{S_{1}}),C(L_{S_{2}})),$ (5)
where unary potential $\psi_{u}(C(L_{S}))$ is defined with negative logarithm of a probability distribution $p_{C(L_{S})}(C)$ over semantic labels as follows:
$\begin{split}&p_{C}(C(L_{S}))=\frac{P_{\nu}^{C(L_{S})}(L_{S})}{\sum_{C}P_{\nu}^{C}(L_{S})},\\ &\psi_{u}(C(L_{S}))=-\log p_{C}(C(L_{S})).\end{split}$ (6)
Binary potential $\psi_{p}(C(L_{S_{1}}),C(L_{S_{2}}))$ is as follows:
@@
$\small\begin{split}&k(L_{S_{1}},L_{S_{2}},C(L_{S_{1}}),C(L_{S_{2}}))=\frac{{P_{\nu}}^{C(L_{S_{1}})}({L_{S_{1}}}){P_{\nu}}^{C(L_{S_{2}})}({L_{S_{2}}})}{\sum_{C(L_{S})}P_{\epsilon}^{C(L_{S})}(L_{S_{1}},L_{S_{2}})},\\ &\psi_{p}(C(L_{S_{1}}),C(L_{S_{2}}))=\mathbb{1}_{C(L_{S_{1}})\neq C(L_{S_{2}})}K_{C}\exp{\left(-\frac{k(*)}{2\theta^{2}}\right)},\end{split}$ (7)
where $k(*)$ is $k(L_{S_{1}},L_{S_{2}},C(L_{S_{1}}),C(L_{S_{2}}))$ and it indicates how likely $L_{S_{1}}$ and $L_{S_{2}}$ stem from different semantic-instance labels. It is straightforward to see that the energy is submodular and semi-metric and, thus, can be efficiently minimized by the $\alpha-\beta$ swap algorithm. Note that to speed up processing, semantic regularization runs only in case of semantic query, i.e. mesh generation or query for planning.
@@
III-D Instance Refinement
This section focuses on improving instance segmentation. Existing methods like [1, 21] commonly face issues of under-segmentation, where two or more instances with the same semantic label get merged. This merging is often caused by under-segmentation in 2D frames or 3D superpoints. Conversely, over-segmentation can occur because of unreliable data association from noisy camera poses and depth measurements. To tackle these challenges, we employ instance refinement on the instance labels. We run refinement separately for each semantic class as the semantic labels were determined previously. This entails running the refinement on the superpoint set associated with a particular semantic label.
Suppose that we are given a semantic label $C$ and the set $\mathcal{V}_{C}$ of superpoints assigned to $C$, and initial instance guess for superpoints $\mathcal{O}_{C}=\{O_{L_{S}}\}$, where ${L_{S}}\in\mathcal{V}_{C}$. The goal is to optimize $\mathcal{O}_{C}$. For each unique instance label $O^{C}_{i}\in\mathcal{O}_{C}$, we define the instance confidence $P_{O}(O^{C}_{i})$, using the graph from the previous section, as the largest edge confidence given the semantic class as follows:

### v2

@@ Abstract
We introduce an online 2D-to-3D semantic instance mapping algorithm aimed at generating comprehensive, accurate, and efficient semantic 3D maps suitable for autonomous agents in unstructured environments. The proposed approach is based on a Voxel-TSDF representation used in recent algorithms. It introduces novel ways of integrating semantic prediction confidence during mapping, producing semantic and instance-consistent 3D regions. Further improvements are achieved by graph optimization-based semantic labeling and instance refinement. The proposed method achieves accuracy superior to the state of the art on public large-scale datasets, improving on a number of widely used metrics. We also highlight a downfall in the evaluation of recent studies: using the ground truth trajectory as input instead of a SLAM-estimated one substantially affects the accuracy, creating a large gap between the reported results and the actual performance on real-world data. The source code will be made public.
Ours
Voxblox++ [1]
Han et al. [2]
INS-CONV [3]
Figure 1: 3D panoptic segmentation comparison: the proposed method versus Voxblox++ [1], Han et al. [4], and INS-CONV [3]. In this example, all competitors under-segment the bookshelf (top rectangle) and over-segment the table (bottom). The proposed technique sets a new benchmark in 2D-to-3D semantic-instance segmentation by ensuring semantic consistency across super-points (a set of adjacent 3D voxels) through graph-based semantic optimization and instance refinement.
@@ I INTRODUCTION
3D semantic-instance mapping is the problem of simultaneously detecting and segmenting objects with their semantic and instance labels to generate a 3D semantic map of the surrounding environment. There are two popular approaches to this problem: 3D-to-3D, which analyzes input 3D point clouds – in this scenario the environment is previously reconstructed; and 2D-to-3D, which analyzes a set of 2D images, further projecting and aggregating predictions on the 3D map – here, reconstruction and semantic mapping is done synchronously. In practice, 3D-to-3D approaches often suffer from a lack of large-scale training data and speed issues, i.e., 3D data is more complex than 2D, which makes processing and analyzing 3D point clouds computationally intensive. 2D-to-3D algorithms leverage state-of-the-art 2D semantic-instance segmentation algorithms hence resolving some of the issues with 3D-to-3D approaches. However, having only 2D observations of the 3D instances, such methods often produce a noisy segmentation of the environment due to the limited perspective of 2D images and the noise in semantic detections per frame and in the camera calibration.
@@ II RELATED WORK
Among the representative techniques are Voxblox++ [1] and PanopticFusion [16]. Voxblox++ [1] employs both semantic instance and geometric segmentation, converting 2D frames into instance-labeled 3D segments. The process involves the incremental registration of geometric segments detected in depth frames onto global 3D counterparts, subsequently integrating semantic and instance labels within these global segments. Our method is inspired by [1], but it offers substantial enhancements, particularly in terms of semantic accuracy. In [16], panoptic segmentation [17] is applied to extract semantic instances for “things” and semantic labels for collective objects (“stuff”), which help improve the quality of semantic-instance segmentation. SceneGraphFusion [18] builds a scene graph with geometric segments and then applies a graph neural network to predict semantic and instance labels. However, SceneGraphFusion utilizes depth images synthetically rendered from a reconstructed 3D mesh, hence using ground truth data. The recent INS-CONV [3] combines the advantages of 2D-3D and 3D-3D procedures. While it builds the 3D model of the environment incrementally from RGB-D frames, it performs segmentation directly in the 3D space. This, however, comes at the cost of requiring 3D ground truth annotations for training.
@@ III SEMANTIC-INSTANCE MAPPING
The proposed incremental 2D-to-3D instance and semantic mapping approach, illustrated in Fig.2, processes a sequence of RGB-D frames alongside poses estimated by a SLAM pipeline [22]. The procedure encompasses a six-step process for each frame, delineated as follows: (i) 2D panoptic-geometric segmentation, detailed in Sec. III-A; (ii) 3D super-point segmentation, described in Sec. III-B; (iii) super-point graph construction and update, elaborated in Sec. III-C. Subsequent steps are executed at the end of the process: (iv) semantic regularization, as discussed in Sec. III-C; (v) instance refinement, outlined in Sec. III-D; and (vi) mesh generation, following the methodology presented in [1].
Preliminaries. Here, we define crucial concepts and notation utilized throughout this manuscript:
  • Panoptic segment $o$: A semantic-instance mask segmented from an RGB image, depicted in Fig.2 III-A.
  • Geometric segment: A convex 3D patch segmented from a depth image, illustrated in Fig.2 III-A.
  • Surface $s$: 3D points derived from a panoptic-geometric segment in an RGB-D frame, as shown in Fig.2 III-B.
  • Super-point $S$: A collection of 3D voxels that represents a whole or part of object instance, labeled as $L_{S}$, and constructed from frame-wise surfaces $s$ (Fig.2 III-B).
  • Super-point graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$: A graph comprising super-points ${S}$ as nodes and edges that model semantic and spatial relationships, as shown in Fig. 2 III-C, III-D.
@@
Panoptic-geometric segmentation, as depicted in Fig. 2 III-A, is applied to each RGB-D frame $F$. This process extracts convex object surfaces based on the assumption from [19]:
@@
The surface of each real-world object is composed of single or multiple neighboring convex surfaces that share the same semantic-instance label.
Initially, panoptic segmentation [23] is performed on the RGB image to produce a set of semantic instance masks. Concurrently, geometric segmentation [19] is applied to the depth image to extract a set of convex segments. The intersections between the panoptic and geometric segmentation masks are utilized to generate a set of surface masks $\mathcal{F}_{s}={s_{i}}$ for each frame $F$. Consequently, each surface mask $s_{i}$ is uniquely associated with a panoptic semantic instance $o_{i}$ and a geometric segment on a per-frame basis, thus mitigating the risk of geometric and semantic under-segmentation.
Each surface $s_{i}$ is attributed with an instance label $o_{i}$, a semantic category $C(o_{i})$, and a panoptic confidence score:
@@
where $C_{\text{Th}}$ and $C_{\text{St}}$ denote the sets of ”thing” and ”stuff” categories, respectively, as defined in [23]. $\text{Score}(o_{i})$ represents the confidence score for the instance $o_{i}$ from [23]. Utilizing the camera pose and the depth image, the point cloud $\text{Pcl}(s_{i})$ corresponding to $s_{i}$ is projected onto a unified Voxel-TSDF structure, which is incrementally constructed.
III-B Semantically Consistent Super-point Segmentation
In the context of volumetric mapping, surfaces are represented by voxels. Consequently, Assumption 1 extends from individual surfaces to encompass super-points:
@@ III-B Semantically Consistent Super-point Segmentation
Each real-world object comprises one or more spatially contiguous super-points, having the same semantic-instance label.
Given Assumptions 1 and 2, we interpret each surface $s_{i}\in\mathcal{F}_{s}$ from Sec. III-A as a partial observation of a super-point $S$ within a global 3D coordinate system. Consequently, super-points are registered or modified with frame-wise surface observations $s_{i}\in\mathcal{F}_{s}$ throughout the sequence of frames $\mathcal{F}$. The correlation between a surface $s_{i}$ and a super-point is established by computing the overlap $\Pi_{s}(L_{S},s_{i})$, which is the count of overlapping voxels between $s_{i}$ and super-point $L_{S}$ in the Voxel-TSDF map. For every point $p_{i}$ in $\text{Pcl}(s_{i})$, the corresponding voxel $V(p_{i})$ is queried for its super-point label $L_{S}$, incrementing the overlap measure as $\Pi_{s}(L_{S}(V(p_{i})),s_{i}):=\Pi_{s}(L_{S}(V(p_{i})),s_{i})+1$. A surface $s_{i}$ is associated with the super-point labeled $L_{S}$ if the overlap surpasses a predefined threshold. In the absence of sufficient overlap, $s_{i}$ receives a new label $L_{S_{\text{new}}}$. Fig. 3 shows that the super-points segmentation meets Assumption 2.
Figure 3: 3D segments colored by super-point labels, depicting objects (e.g., chair, wall) composed of single or multiple super-points. Each super-point is unique to an object.
To validate Assumption 2, we track $\Pi_{V}(L_{S},V_{i})$, the occurrence count of each candidate super-point with label $L_{S}$ in voxel $V_{i}$. The super-point label $L_{S}^{*}$ for voxel $V_{i}$ is determined as follows:
$L_{S}^{*}(V_{i})=\arg\max_{L_{S}}\Pi_{V}(L_{S},V_{i}),$ (2)
indicating that a voxel is allocated to the super-point receiving the majority of votes. This equation is applied when integrating a new surface $s_{i}$ into the global map and querying the super-point label $L_{S}$ in voxel $V(p_{i})$.
To minimize over-segmentation, super-points $S_{1}$ and $S_{2}$ are merged if they are spatially connected and semantically consistent, fulfilling these criteria by being:
  1. 1.
    Spatially connected: $M(L_{S_{1}},L_{S_{2}})>\theta_{\text{merge}}$;
  2. 2.
    Semantically consistent: Semantic categories ${C}(L_{S_{1}})$ and ${C}(L_{S_{2}})$ either match or at least one matches a background category $C_{0}$. Function ${C}$ maps super-points to their semantic labels, as defined in Sec. III-C.
The spatial connection measure $M(L_{S_{1}},L_{S_{2}})$ quantifies the 3D overlap between super-points $S_{1}$ and $S_{2}$, aggregated over frames ${\mathcal{F}}$, as follows:
$\small M(L_{S_{1}},L_{S_{2}})=\sum_{\{\mathcal{F}\}}\sum_{s_{i}\in\mathcal{F}_{s}}\llbracket S_{1}\in L_{N}(s_{i})\wedge S_{2}\in L_{N}(s_{i})\rrbracket,$ (3)
@@
Figure 4: Super-point segmentation without (left; as in [1]) and with (right) semantic consistency checks as proposed in Sec. III-B. Without semantic consistency, noisy geometric segmentations [19] result in merging the super-points of ”wall” and ”bed”. By distinguishing semantic labels as proposed, such incorrect mergers happen less often.
Integrating semantic consistency into super-point segmentation (example in Fig. 4), ensures adherence to Assumption 2, resulting in enhanced segmentation accuracy. This approach improves upon methods such as [1, 2], which primarily consider spatial overlap. The outcome of this step is a refined set of super-points within the RGB-D frame $\mathcal{F}$.
@@ III-C Semantic Segmentation via Super-Point Graph
The subsequent step of estimating super-points for each frame involves translating them into the global coordinate system and associating them with global object instances.
The initial semantic and instance labeling of super-points leverages the frame counting method used in [2, 1] with each frame of the sequence voting for certain semantic and instance labels for super-points observed in that frame. This technique utilizes a 3D overlap-based data association to map frame-wise 2D instances ${o}$ to 3D global instance labels ${O}$. It forms frame-wise associations of super-points and global instances, assigning super-points to the global label with the highest frame count overlap. We refer readers to [2] for an in-depth understanding. This approach, however, exhibits several limitations: (i) By simplistically assigning super-points to the global instance labels with maximum frame counts, it neglects that the noise of panoptic-geometric segmentation could be different across frames due to viewpoint differences; (ii) The method solely relies on individual observations of each super-point without considering the semantic-instance correlations among super-points. Consequently, noisy 2D panoptic-geometric predictions with high frame counts may erroneously influence label assignments, leading to incorrect semantic-instance labeling of super-points, such as over/under-segmentation of instances. An example of this issue is depicted in Fig. 5.
Figure 5: Under-segmentation of the chair and the printer in the corner using Voxblox++ [1], due to inaccuracies in panoptic segmentation. Instances are shown by color.
To address these challenges, we employ a graph-based representation to collectively process frame-wise semantic-instance data and mitigate the impact of noisy 2D predictions by assigning lower panoptic-spatial confidence to such frames. Subsequently, graph-based optimization refines the semantic-instance labels by leveraging the aggregated global panoptic-geometric information within the super-point graph. Specifically, a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is formulated with super-points serving as vertices ($S\in\mathcal{V}$). Each vertex maintains a semantic score $P_{\nu}^{C}(L_{S})$, indicating the likelihood of vertex with label $L_{S}$ being associated with semantic category $C$. To represent panoptic-spatial relationships between super-points, an edge $\epsilon^{C}=(L_{S_{1}},L_{S_{2}})\in\mathcal{E}$ connects two vertices if the panoptic-spatial confidence $P_{\epsilon}^{C}(L_{S_{1}},L_{S_{2}})>0$. This confidence score, $P_{\epsilon}^{C}(L_{S_{1}},L_{S_{2}})$, quantifies the probability of two super-points $L_{S_{1}}$ and $L_{S_{2}}$ sharing the instance label within semantic class $C$ and are spatially contiguous.
Confidence calculation. Here, we detail the computation of confidences $P_{\nu}^{C}$ and $P_{\epsilon}^{C}$. For each frame, rays projected from the camera center through the 2D panoptic masks $o_{i}$ intersect with the 3D super-point map. Voxels intersected by these rays are grouped into sets $\{V_{o_{i}}^{L_{S}}\}$ based on their super-point label $L_{S}$, as determined by Eq. 2. Following this, $P_{\epsilon}^{C}$ and $P_{\nu}$ are updated with information from the current frame.
For the semantic score $P_{\nu}^{C}$, the confidence assigned to a class $C(o_{i})$ for a super-point $S$ is calculated by aggregating the intersected voxels of $S$ with the rays from mask $o_{i}$:
$P_{\nu}^{C}(L_{S})=P_{\nu}^{C}(L_{S})+P_{o_{i}}\cdot N^{S}_{o_{i}},$ (4)
where $N^{S}_{o_{i}}$ is the voxel count in set $V_{o_{i}}^{L_{S}}$. This implies that the more voxels from super-point $L_{S}$ intersect with rays from the panoptic mask $o_{i}$, the higher the likelihood that super-point $L_{S}$ pertains to the semantic category of $o_{i}$.
Panoptic-spatial confidence $P_{\epsilon}^{C}$ is updated for each pair of super-points $L_{S_{1}}$ and $L_{S_{2}}$ whose voxels intersect with rays from the same panoptic mask $o_{i}$. This indicates the likelihood that the two super-points are part of the same object $o_{i}$:
$\small P_{\epsilon}^{C(o_{i})}(L_{S_{1}},L_{S_{2}}):=P_{\epsilon}^{C(o_{i})}(L_{S_{1}},L_{S_{2}})+P_{o_{i}}\cdot P_{s}\cdot\min(N^{S_{1}}_{o_{i}},N^{S_{2}}_{o_{i}}),$
where $P_{o_{i}}$ is the panoptic confidence of $o_{i}$, as defined in Eq. 1, and $P_{s}$ is the spatial confidence based on the inverse distance of $L_{S_{1}}$ and $L_{S_{2}}$. This approach suggests that the greater the panoptic and spatial confidences, the more probable the two super-points observed within the frame belong to the instance $o_{i}$.
Graph optimization. Here, semantic segmentation based on graph optimization is discussed. The goal is to refine the mapping of super-points to semantic labels by constructing and minimizing an energy function $E(C)$, where $C$ represents the semantic labels assigned to super-points. This optimization employs the $\alpha-\beta$ swap algorithm [24].
The energy function incorporates both unary and binary potentials, formulated as:
$\small E(C)=\sum_{L_{S}\in\mathcal{V}}\psi_{u}(C(L_{S}))+\sum_{(L_{S_{1}},L_{S_{2}})\in\mathcal{E}}\psi_{p}(C(L_{S_{1}}),C(L_{S_{2}})).$ (5)
Unary potential $\psi_{u}(C(L_{S}))$ reflects the cost associated with assigning super-point $S$ to semantic category $C(L_{S})$. It is inversely proportional to the probability of $C(L_{S})$:
$\begin{split}&\psi_{u}(C(L_{S}))=-\log\frac{P_{\nu}^{C(L_{S})}(L_{S})}{\sum_{C}P_{\nu}^{C}(L_{S})},\end{split}$ (6)
with $\sum_{C}$ indicating the sum over all semantic labels.
The binary potential, $\psi_{p}(C(L_{S_{1}}),C(L_{S_{2}}))$, represents the cost when two adjacent super-points are assigned different semantic labels. This potential is defined as:
@@
$\small\begin{split}&\psi_{p}(C(L_{S_{1}}),C(L_{S_{2}}))=\mathbb{1}_{C(L_{S_{1}})\neq C(L_{S_{2}})}K_{C}\exp{\left(-\frac{k(*)}{2\theta^{2}}\right)},\\ &k(L_{S_{1}},L_{S_{2}},C)=\frac{{P_{\nu}}^{C(L_{S_{1}})}({L_{S_{1}}}){P_{\nu}}^{C(L_{S_{2}})}({L_{S_{2}}})}{\sum_{C}P_{\epsilon}^{C}(L_{S_{1}},L_{S_{2}})},\end{split}$ (7)
where $k(L_{S_{1}},L_{S_{2}},C)$ evaluates the likelihood of $L_{S_{1}}$ and $L_{S_{2}}$ originating from different semantic categories, based on their panoptic-spatial scores $P_{\epsilon}^{C}(L_{S_{1}},L_{S_{2}})$. A lower $P_{\epsilon}^{C}(L_{S_{1}},L_{S_{2}})$ score implies a higher probability of differing semantic-instance labels, thereby reducing $k(L_{S_{1}},L_{S_{2}},C)$ and, consequently, the cost $\psi_{p}(C(L_{S_{1}}),C(L_{S_{2}}))$.
The energy function $E(C)$ is designed to be submodular and semi-metric, allowing for efficient minimization through the $\alpha-\beta$ swap algorithm $C^{*}=\arg\min_{C}E(C)$. This optimization refines the semantic labels of super-points, enhancing the overall accuracy of the semantic segmentation.
@@
Figure 6: The generated mesh without (left) and with (right) the proposed graph-based semantic optimization (Sec. III-C) and instance label refinement (Sec. III-C).
III-D Instance Segmentation via Super-Point Graph
This section delves into enhancing instance segmentation to address the challenges of under-segmentation and over-segmentation observed in existing methods like [1, 2]. Under-segmentation occurs when super-points from distinct instances are incorrectly merged under a single instance label. Over-segmentation results in a single instance being split across multiple labels due to errors in panoptic-geometric predictions or inaccuracies in data association from camera poses and depth measurements.
To address these issues, we now focus on refining instance labels. Instance label refinement is conducted for each semantic class, allowing for focused adjustment within super-point sets on specific semantic labels. Unlike semantic categories, which are finite, the range of instance labels is potentially infinite, expanding as new instances are detected during the mapping process. Therefore, the optimization strategies applicable to semantic labeling are unsuitable for instance segmentation, necessitating a different approach.
Our method employs rule-based algorithms for assigning instance labels to super-points. We start with a set of super-points $\mathcal{V}_{C}=\{L_{S}\;|\;C(L_{S})=C\}$ semantically linked to class $C$, along with their initial instance assignments $\mathcal{O}_{C}={O_{L_{S}}}$, where ${L_{S}}\in\mathcal{V}_{C}$. The objective is to refine $\mathcal{O}_{C}$ by clustering super-points with strong interconnections (high panoptic-geometric scores $P{\epsilon}^{C}$) under the same instance label and segregating those with weak links.
An iterative process is employed, beginning with defining the instance confidence $P_{O}(O^{C}_{i})$ for each unique instance label $O^{C}_{i}\in\mathcal{O}_{C}$ as the maximum edge confidence within the semantic class $C$: